OpenStudy (anonymous):

The perimeter of a rectangle is 48 feet. If the length measures 16 feet, which is the width? A. 16 feet B. 8 feet C. 10 feet D. 5 feet

OpenStudy (anonymous):

@NotSObright help

OpenStudy (anonymous):

B. 8 feet

OpenStudy (anonymous):

how do I find the answer out though ? what did you do to get to that answer ?

OpenStudy (anonymous):

Let the width = x Perimeter = sum of all sides = 16 + x + 16 + x means 48 = 16 + x +16 + x 48 = 32 + 2x 48-32 = 2x 16 = 2x 16/2 = x 8 = x

OpenStudy (anonymous):

Is it clear now?

OpenStudy (anonymous):

Yeah thanks .